How do I find reference angles? How do I know which quadrant certain angles are in such as 11pi/3 and 7pi/8? Do I convert them to degrees to figure out the exact spot on the circle before finding the reference angle? How do I figure out the reference angle? Thanks so much!
Answer by stanbon(75887)

How do I find reference angles? How do I know which quadrant certain angles are in such as 11pi/3 and 7pi/8? Do I convert them to degrees to figure out the exact spot on the circle before finding the reference angle? How do I figure out the reference angle?
---
You could convert to degrees if it helps you.
If you work the problem in radians you have
to know the 1st Q goes from 0 to pi/2
the 2nd Q goes from pi/2 to pi
the 3rd Q goes from pi to (3/2)pi
the 4th Q goes from (3/2)pi to 2pi.
-----
General rule: The reference angle is the
|acute angle| between the terminal side and
the x-axis.
--------
Ref. angle of (11/3)pi is ?
(11/3)pi = 3pi + (2/3)pi
(11/3)pi is in the 4th Quadrant
with -(1/3)pi as the acute angle.
Reference angle = (1/3)pi.
=============================
Ref. angle of (7/8)pi is ?
Terminal side is 2nd Q.
Acute angle with x-axis = (1/8)pi
=============================
Cheers,
Stan H.
========